OpenStudy (anonymous):

Suppose you want to end up with $5,000 in a bank account after 4 years earning a rate of 3.5% compounding monthly. How much would you have to initially invest?

OpenStudy (anonymous):

u want to end up wid only 5000 bucks? :P Just kidding Use the compound interest formula: Amount = P*(1 + r/100)^t

OpenStudy (anonymous):

What does p and r equal? T=time right?

OpenStudy (anonymous):

p is intiially what u put in. Called principal r is rate t is time, in this case, no of months, as interest is monthly

OpenStudy (anonymous):

Would the answer be 5007?

OpenStudy (anonymous):

no. show me wat u did

OpenStudy (oaktree):

No, the answer is 4346.79117.

OpenStudy (anonymous):

5000*(1+0.035/100)^4

OpenStudy (oaktree):

You should use: \[Final amount = (Principalamount)\times e ^{rt}\]

OpenStudy (oaktree):

So, 5000 = P x e^(0.035)(4), and of course use e=2.718281828...

OpenStudy (oaktree):

The formula you posted is for depreciating and appreciating values. Not bank accounts.

OpenStudy (anonymous):

Isnt that to be used if we have a copmpounding interest?

OpenStudy (oaktree):

Look at the question. It says "3.5% compounded monthly".

OpenStudy (anonymous):

oki so? monthly and yearly doesnt mattter right?